# Enhancing the use of family planning service statistics using a Bayesian modelling approach to inform estimates of modern contraceptive use in low- and middle-income countries

### v1

@@ 6.3 Global findings: Impact of inclusion of EMU in FPET
Figure 4 illustrates the differences in mCPR point estimates, by percentage points, across all countries in the database, categorized by marital status and EMU data type. A positive difference in mCPR indicates an increase when the EMU data is included compared to when only surveys are used. The use of EMU-clients EMUs results in the largest impact overall, observing a maximum increase in mCPR of 3.5 percentage points (pp), 4.2pp and 1.3pp, for AWRA, MWRA and UWRA estimates respectively. FP users largest impact on point estimates of mCPR was 1.2pp, 1.5pp and 0.8pp on AWRA, MWRA and UWRA respectively. In terms of the use of FP visits, the largest change in point estimates was 1.5pp, 1.6pp, and 1.7pp; having the most impact on UWRA mCPR estimates in 2023. All data types show a positive median difference in mCPR estimates for UWRA, suggesting that supplementing the survey model with the EMU data model generally increases mCPR estimates for UWRA. EMU-clients data has the most variation in terms of point estimate differences, ranging from a decrease of 4.2pp to an increase of 3.7pp when considering MWRA estimates for example. The median point estimate difference when considering the use of FP visits is 1.2pp, 1.3pp and 0.8pp for AWRA, MWRA and UWRA, with the plot highlighting that these are the highest median point estimate differences across all data types. In terms of uncertainty with respect to model estimates, measured by the width of credible intervals, the inclusion of EMUs has no substantial effect.
@@ 6.4 Country-level case studies
Country B provide examples of contexts where observation-level uncertainty associated with EMUs is low, as illustrated by narrow error bars, however we don’t have any EMU data before during survey-informed years to indicate EMUs are tracking mCPR well. In Country B, recent EMU trends appear to be increasing quicker than the model-based mCPR estimates, suggesting a recent increase in modern contraceptive use that has yet to be reflected in survey data. Point estimates of mCPR increase by 1.5pp, 1.6pp and 1pp for AWRA, MWRA and UWRA when including EMUs compared to the survey-only model.
Country C presents a context with limited data prior to the most recent survey. However, the single available observation indicates that EMUs were tracking mCPR trends well during the limited, survey-informed period. Observation-level uncertainty associated with EMUs is increasing, as shown by widening the error bars, which reduces the impact of more recent EMU values. That said, including EMU results leads to a larger estimated increase in mCPR in the years since the most recent survey, compared to using surveys alone. Specifically, increases of 3.5pp for AWRA, 4.2pp for MWRA, and 0.7pp for UWRA are observed compared to the survey-only model.
@@ 7 Discussion
When evaluating the impact of EMU inclusion on 2023 mCPR estimates, we observed maximum impacts of 3.5 percentage points for all women of reproductive age, 4.2 percentage points for married women of reproductive age, and 1.3 percentage points for unmarried women of reproductive age with the use of EMU-clients data. Due to country-level variability, the uncertainty associated with EMUs can be substantial, and in some settings, their inclusion has minimal impact, highlighting one of the model’s strengths.

### v2

@@ 6.3 Global findings: Impact of inclusion of EMU in FPET
Figure 4 illustrates the differences in mCPR point estimates, by percentage points, across all countries in the database, categorized by marital status and EMU data type. A positive difference in mCPR indicates an increase when the EMU data is included compared to when only surveys are used. The use of EMU-clients EMUs results in the largest impact overall, observing a maximum increase in mCPR of 3.5 percentage points (pp), 4.2pp and 1.3pp, for AWRA, MWRA and UWRA estimates respectively. FP users largest impact on point estimates of mCPR was 1.2pp, 1.5pp and 0.8pp on AWRA, MWRA and UWRA respectively. In terms of the use of FP visits, the largest change in point estimates was 1.5pp, 1.6pp, and 1.7pp; having the most impact on UWRA mCPR estimates in 2023. All data types show a positive median difference in mCPR estimates for UWRA, suggesting that supplementing the survey model with the EMU data model generally increases mCPR estimates for UWRA. EMU-clients data has the most variation in terms of point estimate differences, ranging from a decrease of 3.7pp to an increase of 4.2pp when considering MWRA estimates for example. The median point estimate difference when considering the use of FP visits is 1.2pp, 1.3pp and 0.8pp for AWRA, MWRA and UWRA, with the plot highlighting that these are the highest median point estimate differences across all data types. In terms of uncertainty with respect to model estimates, measured by the width of credible intervals, the inclusion of EMUs has no substantial effect.
@@ 6.4 Country-level case studies
Country B provide examples of contexts where observation-level uncertainty associated with EMUs is low, as illustrated by narrow error bars, however we don’t have any EMU data before during survey-informed years to indicate EMUs are tracking mCPR well. In Country B, recent EMU trends appear to be increasing quicker than the model-based mCPR estimates, suggesting a recent increase in modern contraceptive use that has yet to be reflected in survey data. Point estimates of mCPR increase by 1.5pp, 1.7pp and 1pp for AWRA, MWRA and UWRA when including EMUs compared to the survey-only model.
Country C presents a context with limited data prior to the most recent survey. However, the single available observation indicates that EMUs were tracking mCPR trends well during the limited, survey-informed period. Observation-level uncertainty associated with EMUs is increasing, as shown by widening the error bars, which reduces the impact of more recent EMU values. That said, including EMU results leads to a larger estimated increase in mCPR in the years since the most recent survey, compared to using surveys alone. Specifically, increases of 1pp for AWRA, 1.1pp for MWRA, and 0.7pp for UWRA are observed compared to the survey-only model.
@@ 7 Discussion
When evaluating the impact of EMU inclusion on 2023 mCPR estimates, we observed maximum impacts of 3.5 percentage points for all women of reproductive age, 4.2 percentage points for married women of reproductive age, and 1.6 percentage points for unmarried women of reproductive age with the use of EMU data. Due to country-level variability, the uncertainty associated with EMUs can be substantial, and in some settings, their inclusion has minimal impact, highlighting one of the model’s strengths.